How does the 2nd derivative calculate? Please show the calculation.

  Find the first and second derivatives of the function f(x) (x2 2)5 10x (x2 + 2)4 f (x) f )10(x22)3( 9x2 +2)
Answer #1

SECOND DERIVATIVE: [f(e)= f() de2 10 (2) dæ d 10 dx 2) 4 d 10 (a (2 + 2) da 4 = 10 (1(2 +2)+4(2 +2) 2 2 da +2)+4( 2(2 +2)) =